# Instanton-Induced Supersymmetry Breaking in Topological Semimetals

### v1

@@ Abstract
Introduction. Supersymmetry (SUSY) attracts significant attention due to its profound potential for extending the Standard Model [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. To explain the absence of SUSY particles in current experiments, the spontaneous SUSY breaking at the low-energy regime is a necessary ingredient in SUSY quantum field theories [6, 5, 7]. To elucidate this core concept, Witten introduced supersymmetric quantum mechanics (SUSY QM) [5], namely the $(0+1)$D SUSY field theories, where the instanton effect acts as a key mechanism of dynamical SUSY breaking. Since then, it was soon clear that SUSY QM was interesting in its own right, and has been developed to understand various solvable potential problems [13, 14, 15, 16]. In this work, we shall develop a new horizon of SUSY quantum mechanics, through revealing the connection between SUSY QM and non-uniform topological semimetals. Before entering our topic, it would be noteworthy that SUSY theories have been realized or found applications in various fields outside of high-energy physics, such as optics, disordered systems, and condensed matters [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41].
@@
In condensed matter experiments, various non-uniform topological semimetals with translational symmetry breaking have been extensively studied. For instance, the translational symmetry may be violated under magnetic fields or strain, as actively being experimentally examined for topological Dirac, Weyl, and nodal-line semimetals [42, 43, 44, 45, 46, 47, 48, 49, 50]. A central issue in these studies is to determine whether a finite critical field or strain is required for gap opening or the destruction of topological zero modes. Despite significant experimental and theoretical efforts, this question remains unresolved. Particularly, due to the lack of translational symmetry, conventional theoretical methods become ambiguous in capturing the topological essence of such non-uniform systems.
@@
In this work, we establish a rigorous theoretical characterization for such non-uniform topological semimetals in terms of SUSY QM. We systematically implement SUSY QM into a wide range of non-uniform topological semimetals possessing chiral symmetry, where broken translational symmetry, e.g., by a magnetic field, is effectively captured by a SUSY potential. Remarkably, we uncover that the dynamical SUSY breaking, mediated by the instanton effect, serves as the mechanism underlying the gap opening or destruction of these semimetallic phases. The magnitude of energy gap is proportional to instanton effect over the SUSY potential valleys, and can be determined by instanton calculations. Based on the instanton mechanism, we propose a straightforward criterion for determining whether an energy gap has been opened without requiring detailed calculations. Specifically, a finite energy gap is opened if and only if there are an even number of zeros in the SUSY potential.
@@
Our theory yields several unexpected results. We find that the Dirac points in graphene become gapped under a magnetic field, even though the gap size is exponentially small. Furthermore, we demonstrate that pairs of Weyl points are annihilated regardless of the strength of magnetic field when oriented in a specific direction, clarifying that there is no threshold for the annihilation to occur [49, 50]. Finally, we find that even a weak magnetic field can destroy the topologically stable nodal-line semimetallic phases. These phenomena all result from dynamical SUSY breaking via instanton effect, as identified by our proposed method.
@@
where $\sigma_{\pm}=1/2(\sigma_{1}\pm i\sigma_{2})$ and $\sigma_{i}^{\prime}s$ with $i=1,2,3$ are Pauli matrices.
@@
where $\{\mathcal{Q},\mathcal{Q}^{\dagger}\}=\mathcal{Q\mathcal{Q}^{\dagger}}+\mathcal{Q}^{\dagger}\mathcal{Q}$. Note that a different font is used for the SUSY Hamiltonian $\mathcal{H}$. Equations (2) and (3) together constitute the algebra of SUSY QM [17], confirming its emergence in the square of (1).
Correspondence between gapped/gapless topological phases and broken/unbroken SUSY. From Eq. (3), eigenvalues $E$ of $H$ are linked to the SUSY eigenvalues $\mathcal{E}$ of $\mathcal{H}$ through the equation $E=\pm\sqrt{\mathcal{E}}$, where SUSY energy $\mathcal{E}$ is always non-negative. Hence, the SUSY ground-state energy $\mathcal{E}_{g}$ is related to the band gap $\Delta$ of the chiral-symmetric Hamiltonian $H$ as
@@
where $v_{i}$ is a real parameter, $k_{i}$ is the momentum in the $i$-th direction, and $g(\tilde{\mathbf{k}})$ is a real function of the rest momenta $\tilde{\mathbf{k}}$ excluding $k_{i}$. Notably, this model is applicable to graphene [Eqs. (12,15)], Weyl [Eq. (18)], and topological nodal-line [Eq. (19)] semimetals.
@@
$v_{i}k_{i}\rightarrow-iv_{i}\partial_{i},\quad g(\tilde{\mathbf{k}})\rightarrow g(x_{i},\tilde{\mathbf{k}}).$ (9)
By treating $g(x_{i},\tilde{\mathbf{k}})$ as the SUSY potential $w(x)$, we arrive at the Witten’s model (6) up to a prefactor $v_{i}$. It is worth noting that the rich variety of $g(\tilde{\mathbf{k}})$ in chiral-symmetric topological materials allows the realization of different types of Witten’s model.
@@
The square of SUSY potential $w_{\text{single}}(\mathsf{x})=-v_{F}/l_{B}^{2}\text{$\mathsf{x}$}$ is ploted in Fig. 1(c). Evidently, there is just one zero (odd) for a single Dirac point. Hence, using Witten’s criteria, the spectrum is not gapped as SUSY is unbroken, in agreement with previous studies. The full spectrum is derived in the SM [51].
Fig. 1: (a) The momentum-space location of Dirac points in graphene. (b) Comparison of spectrum between the tight-binding model (11) and the $k\cdot p$ model for double Dirac points (15). (c) and (d) show the square of SUSY potential for single Dirac points in Eq. (12), and double Dirac points in Eq. (14). Red arrows indicate the zeros of SUSY potential.
@@
where $w_{\text{double}}(\mathsf{x})=\mu_{y}(K^{2}-\text{$\mathsf{x}$}^{2}/l_{B}^{4})$. As shown in Fig. 1(d), the SUSY potential exhibits two zeros (even) in this case, due to the two Dirac points. Hence, using the Witten’s criteria, the spectrum is gapped due to spontaneous SUSY breaking, in stark contrast with previous studies. We emphasize that the two Dirac points are gapped no matter how weak the field is.
@@
$\Delta=\Delta_{0}\sqrt{K\cdot B}\exp\left[-\lambda\frac{K^{3}}{B}\right],$ (17)
where $\Delta_{0}=\sqrt{8ev_{x}\mu_{y}/\hbar\pi}$, and $\lambda=4\hbar\mu_{y}/3ev_{x}$. Evidently, when $B\neq 0$, no matter how small it is, the Dirac points are gapped, aligning with Witten’s criteria. We find that due to the large separation of Dirac points ($K=8.5\,$nm-1), the gap size is exponentially small, e.g., $\Delta=10$ meV for $B=10^{4}$ T.
@@
Pair annihilation of Weyl points under a weak magnetic field. In experiments [49, 50], it was observed that pairs of Weyl points with opposite chiralities are annihilated when a magnetic field is applied perpendicular to their separation direction. Furthermore, these studies suggested that the field strength must exceed a certain threshold to observe the annihilation. Here, we clarify that no such threshold exists for pair annihilation: the annihilation occurs as long as the magnetic field is present, due to dynamical SUSY breaking.
@@
where the separation $K$ can be small in Weyl semimetals [49, 50]. To determine the gap opening upon applying a perpendicular magnetic field, i.e., in the $z$-direction, it suffices to consider the Hamiltonian at $k_{z}=0$, as the system must be gapped elsewhere [55]. This brings us back to the same scenario for the double Dirac points, since the Hamiltonian reduces to Eq. (15) with chiral symmetry. Given the two zeros in the SUSY potential [Fig. 1(d)], the presence of a magnetic field, regardless of its strength, results in pair annihilation due to the gap opening by instanton effect.
The gap size obtained by instanton calculations in Fig. 2(b) is extremely small for field strength below a certain value due to the nature of the exponential function, which is consistent with the experimental observation of a threshold [56]. However, our analytical results of Eq. (17) unambiguously clarify that as long as the magnetic field exists, the gap opens and the annihilation occurs. Our results is further confirmed by numerical calculations based on the lattice model corresponding to (18), as shown in Fig. 2(b).
Fig. 3: (a) Spectrum of the topological nodal-line semimetal at $k_{x}=0$ (left). The noal line is located on the $(k_{y},k_{z})$-plane (right). (b) The squared SUSY potential $w_{k_{z}}^{2}(\mathsf{z})$ for different $k_{z}$ values. Red arrows highlight the zeros.
Furthermore, we propose an alternative method for probing the pair annihilation, since determining the band gap through transport experiments may be challenging. As shown in Fig. 2(c), we compute the local density of states (LDOS) at zero energy under different magnetic fields. At $B=0$, the Fermi-arc surface states are localized at the two boundaries in the $x$-direction. As the magnetic field increases to $B=20$ T, a noticeable reduction in the boundary LDOS is observed, and at $B=40$ T, the localization is completely destroyed. Therefore, the pair annihilation is also detectable by LDOS, which can be observed in STM experiments [57].
Destruction of topological nodal-line semimetallic phase by spontaneous SUSY breaking. Finally, we apply our theory to topological nodal-line semimetals, with a typical $k\cdot p$ model given by [58, 59, 60]
@@
which has the chiral symmetry $S=\sigma_{3}$. Here, $v_{z}$, $t$, and $r$ are model parameters. As shown in Fig. 3(a), the nodal points form a ring with radius $r$.
@@
where $w_{k_{z}}(\mathsf{x})=t\left[r^{2}(k_{z})-\mathsf{x}^{2}/l_{B}^{4}\right]$ and $r^{2}(k_{z})=r^{2}-k_{z}^{2}$. Unlike previous cases, the SUSY potential $w_{k_{z}}(\mathsf{x})$ now depends on $k_{z}$. As shown in Fig. 3(b), within the nodal ring ($|k_{z}|<r$), the SUSY potential exhibits two zeros, whereas no zeros are found outside the ring ($|k_{z}|>r$). Given that SUSY potential features two zeros within the nodal ring region, the spectrum is gapped throughout this region due to dynamical SUSY breaking. The full spectrum is calculated in the SM [51].
Summary and discussion. We have applied SUSY QM, particularly its core concepts, to a wide range of non-uniform topological semimetals possessing chiral symmetry. We have uncovered the underlying mechanism of a novel kind of gap opening in these materials as dynamical SUSY breaking via the instanton effect, which can be straightforwardly determined using our proposed method. While we focus on topological semimetals, the general framework can be extended to other topological materials in chiral symmetry classes, which constitute half of the ten classifications [61]. Our theory could also find applications in systems where chiral-symmetric models serve as effective descriptions, such as the chiral approximation of twisted bilayer graphene [62, 63].
@@ Acknowledgements.
W.B.R. is grateful to Prof. Shaojie Ma for valuable discussions. This work was supported by the Guangdong-Hong Kong Joint Laboratory of Quantum Matter, the Quantum Science Center of Guangdong-Hong Kong-Macau Greater Bay Area, the NSFC/RGC JRS grant (RGC Grant No. N_HKU774/21, NSFC Grant No. 12161160315), the GRF of Hong Kong (Grant Nos. 17310622 and 17303023), and the CRF of Hong Kong (Grant No. C6009-20G). W.B.R. was supported by the RGC Postdoctoral Fellowship (Ref. No. PDFS2223-7S05).

### v2

@@ Abstract
Introduction. Supersymmetry (SUSY) attracts significant attention due to its profound potential for extending the Standard Model [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. To explain the absence of SUSY particles in current experiments, the spontaneous SUSY breaking at the low-energy regime is a necessary ingredient in SUSY quantum field theories [6, 5, 7]. To elucidate this core concept, Witten introduced supersymmetric quantum mechanics (SUSY QM) [5], namely the $(0+1)$D SUSY field theories, where the instanton effect acts as a key mechanism of dynamical SUSY breaking. Since then, it was soon clear that SUSY QM was interesting in its own right, and has been developed to understand various solvable potential problems [13, 14, 15, 16]. In this work, we shall develop a new horizon of SUSY quantum mechanics, through revealing the connection between SUSY QM and non-uniform topological semimetals. Before entering our topic, it would be noteworthy that SUSY theories have been realized or found applications in various fields outside of high-energy physics, such as optics, disordered systems, and condensed matters [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42].
@@
In condensed matter experiments, various non-uniform topological semimetals with translational symmetry breaking have been extensively studied. For instance, the translational symmetry may be violated under magnetic fields or strain, as actively being experimentally examined for topological Dirac, Weyl, and nodal-line semimetals [43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54]. A central issue in these studies is to determine whether a finite critical field or strain is required for gap opening or the destruction of topological zero modes. Despite significant experimental and theoretical efforts, this question remains unresolved. Particularly, due to the lack of translational symmetry, conventional theoretical methods become ambiguous in capturing the topological essence of such non-uniform systems.
@@
In this work, we establish a rigorous theoretical characterization for such non-uniform topological semimetals in terms of SUSY QM. We systematically implement SUSY QM into a wide range of non-uniform topological semimetals possessing chiral symmetry, where broken translational symmetry, e.g., by a magnetic field, is effectively captured by a SUSY potential. Remarkably, we uncover that the dynamical SUSY breaking, mediated by the instanton effect, serves as the mechanism underlying the gap opening or destruction of these semimetallic phases. The magnitude of the energy gap is proportional to instanton effect over the SUSY potential valleys, and can be determined by instanton calculations. Based on the instanton mechanism, we propose a straightforward criterion for determining whether an energy gap has been opened without requiring detailed calculations. Specifically, a finite energy gap is opened if and only if there are an even number of zeros in the SUSY potential.
@@
Our theory yields several unexpected results. We find that the Dirac points in graphene become gapped under a magnetic field, even though the gap size is exponentially small. Furthermore, we demonstrate that pairs of Weyl points are destructed regardless of the strength of magnetic field when oriented in a specific direction, clarifying that there is no threshold for the destruction to occur [53, 54]. Finally, we find that even a weak magnetic field can destroy the topologically stable nodal-line semimetallic phases. These phenomena all result from dynamical SUSY breaking via instanton effect, as identified by our proposed method.
@@
where $\sigma_{\pm}=1/2(\sigma_{1}\pm i\sigma_{2})$ and the $\sigma_{i}^{\prime}s$ with $i=1,2,3$ are the Pauli matrices.Note that $q$ can be a matrix [55].
@@
where $\{\mathcal{Q},\mathcal{Q}^{\dagger}\}=\mathcal{Q\mathcal{Q}^{\dagger}}+\mathcal{Q}^{\dagger}\mathcal{Q}$. Equations (2) and (3) together constitute the algebra of SUSY QM [17], confirming its emergence in the square of (1).
Correspondence between gapped/gapless topological phases and broken/unbroken SUSY. From Eq. (3), eigenvalues $E$ of $H$ are linked to the SUSY eigenvalues $\mathcal{E}$ of $\mathcal{H}$ through the equation $E=\pm\sqrt{\mathcal{E}}$, where the SUSY energy $\mathcal{E}$ is always non-negative. Hence, the SUSY ground-state energy $\mathcal{E}_{g}$ is related to the band gap $\Delta$ of the chiral-symmetric Hamiltonian $H$ as
@@
where $\mathbf{k}\in\mathbb{R}^{d}$ is the momentum in $d$-dimensional space, $k_{i}$ is the component along direction $i$ with $i\in\{1,2,\dots,d\}$, $\tilde{\mathbf{k}}=(k_{1},\dots,k_{i-1},k_{i+1},\dots,k_{d})$ denotes the momentum excluding $k_{i}$, $v_{i}$ is a real parameter, and $g(\tilde{\mathbf{k}})$ is a real function of $\tilde{\mathbf{k}}$. Notably, this model is applicable to graphene [Eqs. (12,15)], Weyl [Eq. (18)], topological nodal-line [Eq. (19)] semimetals. Moreover, in the SM [56], we show that Eq. (8) can be generalized to systems with higher dimensions and systems with non-linear dispersion in $k_{i}$.
@@
$v_{i}k_{i}\rightarrow-iv_{i}\partial_{i},\quad g(\tilde{\mathbf{k}})\rightarrow g(x_{i},\tilde{\mathbf{k}}),$ (9)
where $\partial_{i}=\partial_{x_{i}}$. By treating $g(x_{i},\tilde{\mathbf{k}})$ as the SUSY potential $w(x_{i})$, we arrive at the one-dimensional Witten’s model (6) in the direction $i$ up to a prefactor $v_{i}$. It is worth noting that the rich variety of $g(\tilde{\mathbf{k}})$ in chiral-symmetric topological materials allows the realization of Witten’s models with different types of SUSY potentials.
@@
The square of SUSY potential $w_{\text{single}}(\mathsf{x})=-v_{F}/l_{B}^{2}\text{$\mathsf{x}$}$ is ploted in Fig. 1(c). Evidently, there is just one zero for a single Dirac point. Hence, using Witten’s criteria, the spectrum is not gapped as SUSY is unbroken, in agreement with previous studies. The full spectrum is derived in the SM [56].
Fig. 1: (a) The momentum-space location of Dirac points in graphene. The dashed line indicates the reciprocal lattice unit cell. (b) Comparison of spectrum between the tight-binding model (11) and the $k\cdot p$ model for double Dirac points (15). (c) and (d) show the square of SUSY potential for single Dirac points in Eq. (12), and double Dirac points in Eq. (14). Red arrows indicate the zeros of SUSY potential.
@@
where $w_{\text{double}}(\mathsf{x})=\mu_{y}(K^{2}-\text{$\mathsf{x}$}^{2}/l_{B}^{4})$. As shown in Fig. 1(d), the SUSY potential exhibits two zeros in this case, due to the two Dirac points. Hence, using the Witten’s criteria, the spectrum is gapped due to spontaneous SUSY breaking, in stark contrast with previous studies. We emphasize that the two Dirac points are gapped no matter how weak the field is.
@@
$\Delta=\Delta_{0}\sqrt{\lambda_{0}^{3}K/l_{B}^{2}}\exp{\left(-\frac{4}{3}\lambda_{0}K^{3}l_{B}^{2}\right)},$ (17)
where $\Delta_{0}=\sqrt{8v_{x}\mu_{y}/\pi\lambda_{0}^{3}}$ and $\lambda_{0}=\mu_{y}/v_{x}$. Evidently, when $B\neq 0$, no matter how small, the Dirac points are gapped, aligning with Witten’s criteria. We find that due to the large separation of Dirac points ($K=8.5\,$nm-1), the gap size is exponentially small, e.g., $\Delta=10$ meV for $B=10^{4}\,\text{T}$.
@@
Gap opening in Weyl semimetals under a weak magnetic field. In experiments [53, 54], it was observed that pairs of Weyl points with opposite chiralities are gapped when a magnetic field is applied perpendicular to their separation direction. Furthermore, these studies suggested that the field strength must exceed a certain threshold to observe the gap opening. Here, we clarify that no such threshold exists for gap opening: it occurs as long as the magnetic field is present, due to dynamical SUSY breaking.
@@
where the separation $K$ can be small in Weyl semimetals [53, 54]. To determine the gap opening upon applying a perpendicular magnetic field (i.e., in the $z$-direction), it suffices to consider the Hamiltonian at $k_{z}=0$, as the projected bands for $k_{z}\neq 0$ already possess a gap [60]. This brings us back to the same scenario for the double Dirac points, since the Hamiltonian at $k_{z}=0$ reduces to Eq. (15) with chiral symmetry. Given the two zeros in the SUSY potential [Fig. 1(d)], the presence of a magnetic field, regardless of its strength, results in the gap opening in Weyl semimetals due to instanton effect.
As shown in Fig. 2(b), the gap size is extremely small for field strength below a certain value due to the nature of the exponential function, consistent with the experimental observation of a threshold [61]. However, our analytical instanton calculation of Eq. (17) unambiguously clarify that as long as the magnetic field exists, the gap opens. Our results are further confirmed by numerical calculations based on the continuum model of (18), as shown in Fig. 2(b) [56].
Fig. 3: (a) Spectrum of the topological nodal-line semimetal at $k_{x}=0$ (left). The noal line is located on the $(k_{y},k_{z})$-plane (right). (b) The squared SUSY potential $w_{k_{z}}^{2}(\mathsf{x})$ for different $k_{z}$ values. Red arrows highlight the zeros.
Experimental proposal. We propose scanning tunneling spectroscopy (STS) as a direct method to investigate the gap opening. STS, well recognized for its ability to measure Landau levels in topological semimetals [62, 63], can directly observe the characteristic scaling of Eq. (17) against the magnetic field strength, thereby confirming the instanton-induced SUSY breaking mechanism.
Additionally, the gap opening influences topological properties, manifesting as observable changes. As shown in Fig. 2(c), we compute the local density of states (LDOS) at zero energy under different magnetic fields in Weyl semimetals. At $B=0\,\text{T}$, the Fermi-arc surface states are localized at the two boundaries in the $x$-direction. As the magnetic field increases to $B=20\,\text{T}$, a noticeable reduction in the boundary LDOS is observed, and at $B=40\,\text{T}$, the localization is strongly suppressed. Therefore, the destruction of Weyl points is also detectable by LDOS, which can be observed in STM experiments [64].
Destruction of topological nodal-line semimetallic phase by spontaneous SUSY breaking. Moreover, we apply our theory to topological nodal-line semimetals, with a typical $k\cdot p$ model given by [65, 66, 67]
@@
which has the chiral symmetry $S=\sigma_{3}$. Here, $v_{x}$, $t$, and $r$ are model parameters. As shown in Fig. 3(a), the nodal points form a ring with radius $r$.
@@
where $w_{k_{z}}(\mathsf{x})=t\left[r^{2}(k_{z})-\mathsf{x}^{2}/l_{B}^{4}\right]$ and $r^{2}(k_{z})=r^{2}-k_{z}^{2}$. Unlike previous cases, the SUSY potential $w_{k_{z}}(\mathsf{x})$ now depends on $k_{z}$. As shown in Fig. 3(b), within the nodal ring ($|k_{z}|<r$), the SUSY potential exhibits two zeros, whereas no zeros are found outside the ring ($|k_{z}|>r$). Therefore, using the Witten’s criteria, the spectrum is gapped due to the instanton effect. The calculation of the full spectrum can be found in the SM [56].
Discussions. We have uncovered SUSY QM in a wide range of non-uniform topological semimetals possessing chiral symmetry, and revealed the underlying mechanism of a novel kind of gap opening in these materials as dynamical SUSY breaking via the instanton effect. Besides the rich variety of topological materials with chiral symmetry [68], our theory could also find applications in systems where chiral-symmetric models serve as effective descriptions [69, 70]. In the SM [56], we have further generalized our theoretical framework to include discussions of higher-order SUSY QM [71] and to investigate the higher-dimensional extensions.
Finally, we note that the SUSY uncovered in this work operates within the framework of SUSY QM, exhibiting fundamental SUSY properties such as Bose-Fermi degeneracy and broken/unbroken SUSY in a simple yet experimentally accessible setting. It would be interesting to investigate if SUSY can similarly arise in more complex contexts, such as SUSY quantum field theories pursued by particle physicists, within condensed matter physics.
@@ Acknowledgements.
W.B.R. is grateful to Shaojie Ma, Dawei Zhai, Chenjie Wang, and Bo Fu for valuable discussions. This work was supported by the Guangdong-Hong Kong Joint Laboratory of Quantum Matter, the Quantum Science Center of Guangdong-Hong Kong-Macau Greater Bay Area, the NSFC/RGC JRS grant (RGC Grant No. N_HKU774/21, NSFC Grant No. 12161160315), and the GRF of Hong Kong (Grants No. 17310622, No. 17303023, and No. 17301224). W.B.R. was supported by the RGC Postdoctoral Fellowship (Ref. No. PDFS2223-7S05).